OpenStudy (anonymous):

What is the final balance for the investment? $50,000 for 3 years at 3% compounded annually

OpenStudy (anonymous):

Don't you just multiply them all otgether?

OpenStudy (anonymous):

do u know the answer or no

OpenStudy (jdoe0001):

\(\bf Balance=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ P=principal(\textit{original amount})\to 5,000\\ n=(\textit{number of times per cycle, annually, thus once a year})\to 1\\ r=rate\to 3\%\to \frac{3}{100}\to 0.03\\ t=years\to 3 \\ \quad \\ Balance=5,000\left(1+\frac{0.03}{1}\right)^{1\cdot 3}\)

OpenStudy (anonymous):

still don't get it

OpenStudy (jdoe0001):

hmm you may want to brush up your "compound interest" formula section

OpenStudy (jdoe0001):

hmm actually the original amount is 50,000... anyow.. well... same just plug that in the compound interest formula \(\bf Balance=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ P=principal(\textit{original amount})\to 50,000\\ n=(\textit{number of times per cycle, annually, thus once a year})\to 1\\ r=rate\to 3\%\to \frac{3}{100}\to 0.03\\ t=years\to 3 \\ \quad \\ Balance=50,000\left(1+\frac{0.03}{1}\right)^{1\cdot 3}\)

OpenStudy (anonymous):

so the answer will be

OpenStudy (anonymous):

@jdoe0001 so the answer will what

OpenStudy (jdoe0001):

plug in your values and get the balance

OpenStudy (tkhunny):

@tydynamitenewaccount Your efforts to goad others into doing your work for you and doing your homework for you are not consistent with the intent of this Website. Please read the Code of Conduct and heed its wisdom. Show YOUR work. That is the plan, here.